[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue May 30, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue May 30 19:19:49 UTC 2017


HSU Shift Log for Tue May 30 2017 15:19:43 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: SHIPS-ECMWF did not run properly for EP912017 at 1200 UTC, but  
1200 UTC data was available by 1800 UTC.

N-AWIPS issues: N-AWIPS continues to occasionally not show the "Accept"  
button to display various products.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170530/f5eaffd4/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list